# PoseStreamer: A Multi-modal Framework for 6DoF Pose Estimation of Unseen Moving Objects

### v1

@@ Abstract
Six degree of freedom (6DoF) pose estimation for novel objects is a critical task in computer vision, yet it faces significant challenges in high-speed and low-light scenarios where standard RGB cameras suffer from motion blur. While event cameras offer a promising solution due to their high temporal resolution, current 6DoF pose estimation methods typically yield suboptimal performance in high-speed object moving scenarios. To address this gap, we propose PoseStreamer, a robust multi-modal 6DoF pose estimation framework designed specifically on high-speed moving scenarios. Our approach integrates three core components: an Adaptive Pose Memory Queue that utilizes historical orientation cues for temporal consistency, an Object-centric 2D Tracker that provides strong 2D priors to boost 3D center recall, and a Ray Pose Filter for geometric refinement along camera rays. Furthermore, we introduce MoCapCube6D, a novel multi-modal dataset constructed to benchmark performance under rapid motion. Extensive experiments demonstrate that PoseStreamer not only achieves superior accuracy in high-speed moving scenarios, but also exhibits strong generalizability as a template-free framework for unseen moving objects.
@@ I Introduction
6DoF pose estimation[1] aims to compute the rigid six-dimensional transformation (3D translation and rotation) between an object and a camera. Traditional 6DoF pose estimation methods[2, 3, 4] cannot be directly applied to unseen novel objects during the training phase. In practical scenarios, different applications often provide diverse input objects. More recent efforts have focused on real-time pose estimation for arbitrary novel objects. However, existing RGB camera-based methods[5] are prone to motion blur, which particularly limits their performance in high-speed low-lit scenarios.
@@
Figure 1: Overview of the PoseStreamer Architecture. The framework proceeds in three stages: (A) Unseen object initialization via RGB-based reconstruction and the Adaptive Pose Memory Queue (AMQ). (B) High-speed 3D center estimation via the Multi-modality 3D Tracker (M3D) on SD and TD streams. (C) Fine-grained 6DoF optimization via the Ray Pose Filter (RPF). The filter samples and selects pose hypotheses along the camera ray.
@@
Some studies introduce event cameras[6] which capture motion information via luminance variations. It outputs a sparse event stream with high temporal resolution. Such a stream can act as a complementary modality to the motion geometry information for RGB cameras. Such a stream can act as a complementary modality to the motion geometry information for RGB cameras. Currently, relevant research on visual object tracking using event cameras remains relatively limited. EventVOT [7] distills knowledge from an RGB-Event dual-modal teacher network into a pure event-driven student network. AFNet [8] fuses frame-based RGB data and event-based temporal data through a multi-modal alignment and fusion module. STNet [9] integrates Transformer with spiking neural networks to build an event-driven tracking framework. However, all the aforementioned methods only validate 2D tracking performance and do not explore object tracking in the 3D space. Our research aims to leverage the RGB-Event fused modality to achieve accurate 6DoF pose estimation for arbitrary moving objects. On the one hand, event cameras only capture edge information from luminance variations. They lack critical contextual details, such as texture, color, and global semantics, which are essential for object orientation estimation. Moreover, a significant distribution gap exists between event data and RGB modality data[10, 11], posing a barrier to direct fusion for information complementarity. On the other hand, event cameras output relatively sparse spatial signals. This sparsity can lead to tracking failure when target object contours are insufficiently distinct.[12] In addition, the temporal resolution mismatch between event and RGB cameras further amplifies the uncertainty of spatial perception.
@@
To address the aforementioned challenges, this paper proposes PoseStreamer, a multi-modal 6DoF pose estimation framework tailored for high-speed moving scenarios. Its core architecture comprises three key components: (1) The Adaptive Pose Memory Queue (AMQ), a sliding-window module that leverages historical orientation cues to guide current-frame orientation inference. (2) The Multi-modality 3D tracker (M3D), which supplies robust 2D priors to enhance the recall of 3D object centers. (3) The Ray Pose Filter (RPF), designed to refine pose estimates along camera ray directions. Furthermore, to evaluate performance under high-speed motion, we introduce MoCapCube6D, a novel multi-modal pose estimation dataset that facilitates comprehensive benchmarking of our method against state-of-the-art approaches. Notably, the proposed method exhibits strong generalizability and can be extended to a template-free framework for novel object instances. Our contributions are summarized as follows:
@@
  • We demonstrate the strong generalizability of our approach by extending it to a template-free framework capable of handling arbitrary novel object instances without retraining, addressing the limitations of traditional methods on unseen objects.
@@ II Related Works
6DoF pose estimation aims to infer the 3D translation and rotation of a target object. Traditional methods often require instance- or category-level CAD models for offline training or template matching [2], which restricts their application to novel unknown objects. Although recent generalizable works [13] relax these assumptions, they typically rely on pre-captured reference views of the test object. In the tracking domain, methods like BundleTrack [14] attempt to generalize to novel objects instantly without prior templates. However, these RGB-based approaches rely heavily on clear textures and accurate feature matching. Consequently, they suffer significant performance degradation in high-speed scenarios where severe motion blur[15] and artifacts sever the required 2D-3D correspondences.
@@ 19 $\hat{R}\leftarrow\mathcal{E}(\hat{R}^{\prime})$;
III-B SD & TD 3D Center Tracking
Conventional pose estimators relying solely on RGB frames often falter in high-speed scenarios due to severe motion blur and large inter-frame displacements. To overcome this limitation, we leverage complementary modalities: Temporal Difference (TD), which captures high-frequency motion changes, and Spatial Difference (SD), which highlights structural edges. Integrating these inputs, we propose M3D, a lightweight and modality-agnostic tracker. As illustrated in Fig. 2, the pipeline extracts features from these stereo streams (RGB, TD, or SD), clusters them based on motion consistency, and robustly computes the 3D center $\mathbf{C}$ via triangulation.
@@ III-B SD & TD 3D Center Tracking
Figure 2: Details of Multi-modality 3D tracker. Features $F_{L}$ and $F_{R}$ are extracted from the left and right Tianmouc cameras, respectively, with modalities including RGB, TD and SD. Feature points are clustered into groups according to motion consistency, after which the 2D centroids $C_{2D}^{L}$ and $C_{2D}^{R}$ are computed from the left and right views. Finally, the 3D object center $\mathbf{C}$ is obtained via disparity-based stereo triangulation.
@@ III-C3 Pose Refinement
where $\tilde{C}$ and $\tilde{R}$ are the final outputs enqueued into the AMQ for the next iteration. Thanks to this data-driven design[20], the refinement decoder is universally applicable to standard RGB, event, and SD inputs.
@@ IV-A1 MoCapCube6D
We present MoCapCube6D, a high-precision benchmark for high-speed tracking featuring a Rubik’s Cube with synchronized RGBTD and SD streams. Ground truth is obtained via a calibrated MoCap system, ensuring sub-millimeter translation and sub-degree rotation accuracy. The dataset includes three scenes with distinct motion patterns (details in Tab. I) and is further partitioned by projected pixel velocity $v$ into Regular ($v<45$ px/s), Medium ($45\leq v<180$), and Faster ($v\geq 180$) to evaluate robustness under diverse dynamics.
@@ IV-A2 YCB Object Set
We also adopt the standard YCB object set[21]. Due to the lack of high-speed SD/TD benchmarks, we collected and labeled real-world sequences of YCB objects using the aforementioned MoCap system. This enables comprehensive qualitative and quantitative validation of our template-free framework on unseen objects.
@@ V Conclusion
In this paper, we present PoseStreamer, a robust multi-modal framework designed for 6-DoF pose estimation of unseen moving objects in high-speed scenarios. To overcome the limitations of standard RGB cameras under rapid motion, we introduce three core components: the Multi-modality 3D Tracker (M3D), which leverages stereo-based multi-modal features to provide reliable 3D center priors. The Adaptive Pose Memory Queue (AMQ), which ensures temporal orientation consistency by utilizing historical cues. And the Ray Pose Filter (RPF), which effectively mitigates depth uncertainty through geometric refinement along camera rays. Furthermore, we construct MoCapCube6D, a high-precision multi-modal benchmark dataset containing synchronized RGB, TD, and SD streams, to evaluate performance across varying speed profiles. Extensive experiments demonstrate that PoseStreamer significantly outperforms state-of-the-art methods in high-speed settings and exhibits strong generalizability as a template-free framework for novel objects.

### v2

@@ Abstract
Six degree of freedom (6DoF) pose estimation for unseen objects is a critical task in computer vision, yet it faces significant challenges in high-speed and low-light scenarios where standard RGB cameras suffer from motion blur. While event cameras offer a promising solution due to their high temporal resolution, current 6DoF pose estimation methods typically yield suboptimal performance in high-speed object moving scenarios. To address this gap, we propose PoseStreamer, a robust multi-modal 6DoF pose estimation framework designed specifically on high-speed moving scenarios. Our approach integrates three core components: an Adaptive Pose Memory Queue that utilizes historical orientation cues for temporal consistency, an Object-centric 2D Tracker that provides strong 2D priors to boost 3D center recall, and a Ray Pose Filter for geometric refinement along camera rays. Furthermore, we introduce MoCapCube6D, a novel multi-modal dataset constructed to benchmark performance under rapid motion. Extensive experiments demonstrate that PoseStreamer not only achieves superior accuracy in high-speed moving scenarios, but also exhibits strong generalizability as a template-free framework for unseen moving objects.
@@ I Introduction
6DoF pose estimation[1] aims to compute the rigid six-dimensional transformation (3D translation and rotation) between an object and a camera. Traditional 6DoF pose estimation methods[2, 3, 4] cannot be directly applied to unseen objects during the training phase. In practical scenarios, different applications often provide diverse input objects. More recent efforts have focused on real-time pose estimation for arbitrary unseen objects. However, existing RGB camera-based methods[5] are prone to motion blur, which particularly limits their performance in high-speed low-lit scenarios.
@@
Figure 1: Overview of the PoseStreamer Architecture. The framework proceeds in three stages: (A) Unseen object initialization via RGB-based reconstruction and the Adaptive Pose Memory Queue (AMQ). (B) High-speed 3D center estimation via the Multi-modality 3D Tracker (M3D) on RGB images and Event streams. (C) Fine-grained 6DoF optimization via the Ray Pose Filter (RPF). The filter samples and selects pose hypotheses along the camera ray.
@@
Some studies introduce event cameras[6] which capture motion information via luminance variations. It outputs a sparse event stream with high temporal resolution. Such a stream can act as a complementary modality to the motion geometry information for RGB cameras. Currently, relevant research on visual object tracking using event cameras remains relatively limited. EventVOT [7] distills knowledge from an RGB-Event dual-modal teacher network into a pure event-driven student network. AFNet [8] fuses frame-based RGB data and event-based temporal data through a multi-modal alignment and fusion module. STNet [9] integrates Transformer with spiking neural networks to build an event-driven tracking framework. However, all the aforementioned methods only validate 2D tracking performance and do not explore object tracking in the 3D space. Our research aims to leverage the RGB-Event fused modality to achieve accurate 6DoF pose estimation for arbitrary moving objects. On the one hand, event cameras only capture edge information from luminance variations. They lack critical contextual details, such as texture, color, and global semantics, which are essential for object orientation estimation. Moreover, a significant distribution gap exists between event data and RGB modality data[10, 11], posing a barrier to direct fusion for information complementarity. On the other hand, event cameras output relatively sparse spatial signals. This sparsity can lead to tracking failure when target object contours are insufficiently distinct.[12] In addition, the temporal resolution mismatch between event and RGB cameras further amplifies the uncertainty of spatial perception.
@@
To address the aforementioned challenges, this paper proposes PoseStreamer, a multi-modal 6DoF pose estimation framework tailored for high-speed moving scenarios. Its core architecture comprises three key components: (1) The Adaptive Pose Memory Queue (AMQ), a sliding-window module that leverages historical orientation cues to guide current-frame orientation inference. (2) The Multi-modality 3D tracker (M3D), which supplies robust 2D priors to enhance the recall of 3D object centers. (3) The Ray Pose Filter (RPF), designed to refine pose estimates along camera ray directions. Furthermore, to evaluate performance under high-speed motion, we introduce MoCapCube6D, a novel multi-modal pose estimation dataset that facilitates comprehensive benchmarking of our method against state-of-the-art approaches. Notably, the proposed method exhibits strong generalizability and can be extended to a template-free framework for unseen object instances. Our contributions are summarized as follows:
@@
  • We demonstrate the strong generalizability of our approach by extending it to a template-free framework capable of handling arbitrary unseen object instances without retraining, addressing the limitations of traditional methods on unseen objects.
@@ II Related Works
6DoF pose estimation aims to infer the 3D translation and rotation of a target object. Traditional methods often require instance- or category-level CAD models for offline training or template matching [2], which restricts their application to unseen objects. Although recent generalizable works [13] relax these assumptions, they typically rely on pre-captured reference views of the test object. In the tracking domain, methods like BundleTrack [14] attempt to generalize to unseen objects instantly without prior templates. However, these RGB-based approaches rely heavily on clear textures and accurate feature matching. Consequently, they suffer significant performance degradation in high-speed scenarios where severe motion blur[15] and artifacts sever the required 2D-3D correspondences.
@@ 19 $\hat{R}\leftarrow\mathcal{E}(\hat{R}^{\prime})$;
III-B Multi-modal 3D Center Tracking
Conventional pose estimators relying solely on RGB frames often falter in high-speed scenarios due to severe motion blur and large inter-frame displacements. To overcome this limitation, we leverage complementary modalities: RGB images, which captures high-frequency motion changes, and Event stream, which highlights structural edges. Integrating these inputs, we propose M3D, a lightweight and modality-agnostic tracker. As illustrated in Fig. 2, the pipeline extracts features from these stereo streams (RGB and event steam), clusters them based on motion consistency, and robustly computes the 3D center $\mathbf{C}$ via triangulation.
@@ III-B Multi-modal 3D Center Tracking
Figure 2: Details of Multi-modality 3D tracker. Features $F_{L}$ and $F_{R}$ are extracted from the left and right cameras, respectively, with modalities including RGB images and event stream. Feature points are clustered into groups according to motion consistency, after which the 2D centroids $C_{2D}^{L}$ and $C_{2D}^{R}$ are computed from the left and right views. Finally, the 3D object center $\mathbf{C}$ is obtained via disparity-based stereo triangulation.
@@ III-C3 Pose Refinement
where $\tilde{C}$ and $\tilde{R}$ are the final outputs enqueued into the AMQ for the next iteration. Thanks to this data-driven design[20], the refinement decoder is universally applicable to standard RGB and event stream inputs.
@@ IV-A1 MoCapCube6D
We present MoCapCube6D, a high-precision benchmark for high-speed tracking featuring a Rubik’s Cube with synchronized RGB and event streams. Ground truth is obtained via a calibrated MoCap system, ensuring sub-millimeter translation and sub-degree rotation accuracy. The dataset includes three scenes with distinct motion patterns (details in Tab. I) and is further partitioned by projected pixel velocity $v$ into Regular ($v<45$ px/s), Medium ($45\leq v<180$), and Faster ($v\geq 180$) to evaluate robustness under diverse dynamics.
@@ IV-A2 YCB Object Set
We also adopt the standard YCB object set[21]. Due to the lack of high-speed multi-modal benchmarks, we collected and labeled real-world sequences of YCB objects using the aforementioned MoCap system. This enables comprehensive qualitative and quantitative validation of our template-free framework on unseen objects.
@@ V Conclusion
In this paper, we present PoseStreamer, a robust multi-modal framework designed for 6-DoF pose estimation of unseen moving objects in high-speed scenarios. To overcome the limitations of standard RGB cameras under rapid motion, we introduce three core components: the Multi-modality 3D Tracker (M3D), which leverages stereo-based multi-modal features to provide reliable 3D center priors. The Adaptive Pose Memory Queue (AMQ), which ensures temporal orientation consistency by utilizing historical cues. And the Ray Pose Filter (RPF), which effectively mitigates depth uncertainty through geometric refinement along camera rays. Furthermore, we construct MoCapCube6D, a high-precision multi-modal benchmark dataset containing synchronized RGB and event streams, to evaluate performance across varying speed profiles. Extensive experiments demonstrate that PoseStreamer significantly outperforms state-of-the-art methods in high-speed settings and exhibits strong generalizability as a template-free framework for unseen objects.